3(x+y)+4(x+y)-6 when x=4 andy=-1
Answer by BurningFlame_marc(77)

Hello!
All you need to do is substitute x and y with 4 and 1 respectively.
Solution:
3(x+y)+4(x+y)-6
=3(4+-1)+4(4+-1)-6
=12-3+16-4-6
=9+12-6
=9+6
=15
Final Answer: 15